The width of a rectangle is 20 cm. The diagonal is 8 cm more than the
length. Find the length of the rectangle.
a.

Answer by ewatrrr(24785)

Applying the Pythagorean theorem
(L+8)^2 - 400 = L^2
L^2 + 16L + 64 - 400 = L^2
16L + 64 = 400
16L = 336
L = 21cm